# COMPASS: Robust Feature Conformal Prediction for Medical Segmentation Metrics

### v1

@@ Abstract
In clinical applications, the utility of segmentation models is often based on the accuracy of derived downstream metrics such as organ size, rather than by the pixel-level accuracy of the segmentation masks themselves. Thus, uncertainty quantification for such metrics is crucial for decision-making. Conformal prediction (CP) is a popular framework to derive such principled uncertainty guarantees, but applying CP naively to the final scalar metric is inefficient because it treats the complex, non-linear segmentation-to-metric pipeline as a black box. We introduce COMPASS, a practical framework that generates efficient, metric-based CP intervals for image segmentation models by leveraging the inductive biases of their underlying deep neural networks. COMPASS performs calibration directly in the model’s representation space by perturbing intermediate features along low-dimensional subspaces maximally sensitive to the target metric. We prove that COMPASS achieves valid marginal coverage under exchangeability and nestedness assumptions. Empirically, we demonstrate that COMPASS produces significantly tighter intervals than traditional CP baselines on four medical image segmentation tasks for area estimation of skin lesions and anatomical structures. Furthermore, we show that leveraging learned internal features to estimate importance weights allows COMPASS to also recover target coverage under covariate shifts. COMPASS paves the way for practical, metric-based uncertainty quantification for medical image segmentation.
@@ 1 Introduction
Uncertainty quantification is of critical need in medical image analysis, a field used for decision support in high-stakes clinical diagnosis and treatment planning applications [1, 2]. A fundamental task in medical image analysis is image segmentation, the task of separating anatomical structures and lesions from each other within an image. Deep learning models, particularly U-Net variants [3, 4], have achieved state-of-the-art performance in medical image segmentation. In practice, the outputs of these models (“segmentation maps”) are often treated as an intermediate result that are then used to automatically derive downstream metrics of interest (known as “radiomics”), such as the areas/volumes or texture patterns of specific anatomic regions (Figure 1, left). These derived metrics are then used for decision support to guide clinicians in diagnosis and treatment.
@@
Conformal prediction (CP) has emerged as a popular, statistically principled uncertainty quantification framework of choice in machine learning, providing guarantees without restrictive distributional assumptions [5, 6, 7, 8, 9]. While well-studied in the context of typical prediction tasks involving scalar output variables, CP is less explored for tasks such as medical image segmentation, in which the output variables are images. Existing CP methods for segmentation typically focus on deriving bounds for pixel-level errors  [10, 11, 12, 13], which, while useful for understanding variations of local segmentation contours, may yield meaningless or misaligned intervals for downstream derived metrics. On the other hand, a recent study shows that treating the segmentation-to-metric pipeline as a black box and performing CP directly on the output metric space yields intervals that are well-aligned to the metrics (by construction), but are also often inefficient (i.e., large) because the internal biases of the pipeline are not exploited in the vanilla CP formulation [14].
@@
To achieve more efficient intervals, a promising direction is to leverage the powerful inductive biases of neural networks by performing CP on their intermediate representations. Feature Conformal Prediction (FCP) [15, 16, 17] shows that by working in a semantic feature space, it is possible to generate provably tighter prediction intervals. However, the FCP algorithm requires solving a complex optimization to find the closest adversarial feature vector for each data point , which is computationally prohibitive for high-dimensional feature spaces of typical CNN and transformer architectures which are the workhorses of modern medical image segmentation.
@@
To bridge this gap, we introduce Conformal Metric Perturbation Along Sensitive Subspaces (COMPASS), a framework to perform feature CP in a tractable manner to generate valid and efficient prediction intervals for any differentiable metric derived from the output of a neural network. The core concept behind COMPASS is to linearly perturb the features outputted by the network at a particular layer along data-specific directions that are highly sensitive to the metric of interest (Figure 1, center). To make this process tractable for typical neural network layers with a large number of features, we propose computing a low-dimensional manifold of any given layer by applying principal component analysis (PCA) on the gradients of the output metric with respect to each of the layer’s features. This manifold represents the principal directions of sensitivity of the output with respect to that layer. We prove that linear perturbations in the latent space achieve marginal coverage in the output metric space under exchangeability and nestedness assumptions, i.e., for a fresh test point, the ground truth will be contained within the prediction intervals generated by linear feature perturbations. Furthermore, we show that a simple weighted variant of COMPASS may be used to correct for covariate shifts.
@@
We evaluated COMPASS on public datasets for four medical image segmentation tasks: colorectal cancer in histopathology images (EBHI)  [18], skin lesion segmentation (HAM10000) [19], thyroid nodule segmentation in ultrasound images (TN3K) [20], and polyp segmentation in endoscopic images (Kvasir) [21]. Results show that while standard CP methods achieve valid coverage, they often produce unnecessarily wide prediction intervals. On the other hand, COMPASS finds semantically meaningful directions in the latent space that correspond to monotonic metric changes, resulting in efficient intervals. Furthermore, the weighted extension of COMPASS recovers target coverage under covariate shifts and is the most efficient across all weighted baseline methods. COMPASS paves the way for practical, metric-based uncertainty quantification for medical image segmentation.
@@ 2 Method
Let $(X,Y)$ denote an input image and its corresponding ground-truth metric randomly drawn from an unknown distribution $D$. We consider a segmentation model that we can conceptually decompose into a feature extractor $f$ and a subsequent downstream predictor $g$ (see Figure 1-left). The function $f$ maps the input image to a representation, $\hat{Z}=f(X)$, which can be the output of any intermediate layer of the network. The function $g$ then maps this representation to the segmentation mask. Finally, a differentiable metric function $h$ maps the final mask to a scalar prediction, $)\hat{Y}=h(g(\hat{Z}$)). We use lowercase letters to denote the specific values of a particular realization $(x,y)$: $\hat{z}=f(x)$ and $)\hat{y}=h(g(\hat{z}$)).
Figure 1: Overview of COMPASS. (Left) A medical image segmentation network predicts a segmentation “map” from an input image. We conceptually decompose this network into a function $f$ which maps the image to latent features $\hat{z}$, and a function $g$ that maps $\hat{z}$ to the output map. The map may then be used to compute a (differentiable) downstream metric $\hat{y}$ via the function $h$. (Center) We linearly perturb calibration features $\hat{z_{i}}$ in a sample-specific direction $\Delta_{i}$ to find the scores $R_{i}$. The scores are used to find the conformal quantile $\hat{\beta}$. (Right) At test time for subject $n+1$, we perturb the features $\hat{z}_{n+1}$ in the direction $\Delta_{n+1}$ with magnitude $\hat{\beta}$. Under the assumptions of nestedness and exchangeability, the resulting prediction interval achieves marginal coverage (bottom).
@@ 2.1 Theoretical Coverage under Linear Latent Perturbations
To provide rigorous uncertainty quantification for the metric $Y$, we consider linear perturbations in latent space along a direction $\Delta\in\mathcal{Z}$ and define intervals in the metric space. We first formalize the nestedness condition, which is a fundamental requirement for the validity of any conformal procedure [5, 6]. This condition guarantees that a larger perturbation magnitude yields a larger (or equal-sized) prediction set, which is necessary for the quantile-based coverage proof to hold. We explicitly state this assumption because, while trivially satisfied by construction in standard output-space methods, nestedness becomes a non-trivial condition in deep feature spaces with non-linear mappings.
@@
Let $m_{x}:\mathbb{R}\to\mathbb{R}$ denote a real-valued function associated with each input $x\in\mathcal{X}$. For $\beta\geq 0$, define the interval
$S_{\beta}(x):=[\,m_{x}(-\beta),\;m_{x}(+\beta)\,].$
The family $\{S_{\beta}(x)\}_{\beta\geq 0}$ is said to be nested for $x$ if
$\beta_{1}\leq\beta_{2}\;\Rightarrow\;S_{\beta_{1}}(x)\subseteq S_{\beta_{2}}(x).$
@@
Let $(X_{i},Y_{i})_{i\geq 1}$ be exchangeable random pairs with $X_{i}\in\mathcal{X}$ and $Y_{i}\in\mathbb{R}$, and split the data into a training set $D_{\mathrm{tr}}$ and a calibration set $D_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. Using $D_{\mathrm{tr}}$, fit a segmentation model with decoder $g:\mathcal{Z}\to\mathcal{S}$, and let $\hat{z}(x)\in\mathcal{Z}$ denote the latent vector computed deterministically by the trained encoder for input $x$. Let $h:\mathcal{S}\to\mathbb{R}$ be a measurable, differentiable metric, and let $\Delta\in\mathcal{Z}$ be any measurable direction that depends only on $D_{\mathrm{tr}}$.
For $x\in\mathcal{X}$ and $\beta\geq 0$, define
$m_{x}(\beta):=(h\circ g)(\hat{z}(x)+\beta\Delta),\quad S_{\beta}(x):=[m_{x}(-\beta),m_{x}(+\beta)].$
Assume that for every $x$, the family $\{S_{\beta}(x)\}_{\beta\geq 0}$ is nested. For each calibration pair $(X_{i},Y_{i})$, define the non-conformity score
$R_{i}:=\inf\{\beta\geq 0:Y_{i}\in S_{\beta}(X_{i})\}\in[0,\infty],$
@@
$\mathbb{P}\bigl(Y_{n+1}\in S_{\hat{\beta}}(X_{n+1})\;\big|\;D_{\mathrm{tr}}\bigr)\geq 1-\alpha.$
Proof sketch. Exchangeability of $(X_{i},Y_{i})$ ensures that the rank of the test score $R_{n+1}$ among $\{R_{1},\dots,R_{n},R_{n+1}\}$ is uniformly distributed; by construction of $\hat{\beta}$ and the nestedness assumption, the conformal guarantee holds. See Appendix A for full proof.
@@
Theorem 1 can be understood as an instance of conformal risk control [13], where the risk being controlled is the mis-coverage loss for a downstream metric, and the intervals generated by our perturbation scheme form the required nested family of prediction sets. Nestedness ensures that the sets $S_{\beta}(x)$ are weakly monotone with $\beta$. This property guarantees that increasing $\beta$ cannot exclude a previously covered label $Y$, so $R_{i}$ remains the minimal perturbation needed to include $Y_{i}$, analogous to residuals in standard CP. By exchangeability, the test score $R_{n+1}$ is equally likely to occupy any rank among the calibration scores, and the quantile construction of $\hat{\beta}$ ensures the $1-\alpha$ coverage.
@@ 2.2 COMPASS: Conformal Metric Perturbation Along Sensitive Subspaces
The proposed framework COMPASS is built on the insight that we can directly calibrate a downstream metric by perturbing a model’s representations along specific, data-dependent directions, $\Delta$. As our theoretical guarantee (Theorem 1) applies to any measurable perturbation direction, the key to an efficient method lies in choosing a direction that is highly sensitive to the metric of interest. We explore two natural choices for this representation space: the model’s final output logits and a deeper internal feature layer.
The first and simplest approach (COMPASS-L) operates on the model’s final logits. A uniform, scalar shift applied to this tensor directly modulates the model’s overall confidence before the final activation function. This is equivalent to defining the sensitive direction, $\Delta_{i}$, as a tensor of ones. A richer alternative is to perturb an internal feature representation $\hat{z}$, as done in FCP [15]. However, a naive search for an optimal perturbation in high-dimensional spaces is computationally intractable and sensitive to poor local minima. Thus, perturbation directions will not lead to monotonically increasing or decreasing output metrics. To overcome this, we propose a data-driven method to identify a low-dimensional sensitive subspace that is globally effective. We first compute Jacobians of the metric $\hat{y}$ with respect to $\hat{z}$ to provide a local, linear map of the metric’s sensitivity with respect to the features for each training sample:
$J_{i}:=\frac{d\,h(g(\hat{z}_{i}))}{d\hat{z}_{i}}\in\mathbb{R}^{C\times H\times W},$
where $\hat{z}_{i}=f(x_{i})$. Because the full spatial Jacobian is often too high-dimensional to be practical, we apply Principal Component Analysis (PCA) to the set of these vectors from the training set, $\{\mathcal{J}_{i}\}_{i\in D_{\mathrm{tr}}}$, and select the matrix of the top $L$ eigenvectors $V_{L}\in\mathbb{R}^{C\times L}$. In the case of CNNs with image-shaped features, we first sum features per channel before applying PCA.
For any given sample, we find its particular sensitive direction $\mathbf{d}_{i}$ by projecting its sensitivity vector onto this learned subspace, and normalization it to produce direction vector $\Delta_{i}$:

### v2

@@ Abstract
In clinical applications, the utility of segmentation models is often based on the accuracy of derived downstream metrics such as organ size, rather than by the pixel-level accuracy of the segmentation masks themselves. Thus, uncertainty quantification for such metrics is crucial for decision-making. Conformal prediction (CP) is a popular framework to derive such principled uncertainty guarantees, but applying CP naively to the final scalar metric is inefficient because it treats the complex, non-linear segmentation-to-metric pipeline as a black box. We introduce COMPASS, a practical framework that generates efficient, metric-based CP intervals for image segmentation models by leveraging the inductive biases of their underlying deep neural networks. COMPASS performs calibration directly in the model’s representation space by perturbing intermediate features along low-dimensional subspaces maximally sensitive to the target metric. We prove that COMPASS achieves valid marginal coverage under the assumption of exchangeability. Empirically, we demonstrate that COMPASS produces significantly tighter intervals than traditional CP baselines on four medical image segmentation tasks for area estimation of skin lesions and anatomical structures. Furthermore, we show that leveraging learned internal features to estimate importance weights allows COMPASS to also recover target coverage under covariate shifts. COMPASS paves the way for practical, metric-based uncertainty quantification for medical image segmentation.
@@ 1 Introduction
Uncertainty quantification is of critical need in medical image analysis, a field used for decision support in high-stakes clinical diagnosis and treatment planning applications [Begoli et al., 2019, Abdar et al., 2021]. A fundamental task in medical image analysis is image segmentation, the task of separating anatomical structures and lesions from each other within an image. Deep learning models, particularly U-Net variants [Ronneberger et al., 2015, Isensee et al., 2021], have achieved state-of-the-art performance in medical image segmentation. In practice, the outputs of these models (“segmentation maps”) are often treated as an intermediate result that are then used to automatically derive downstream metrics of interest (known as “radiomics”), such as the areas/volumes or texture patterns of specific anatomic regions (Figure 1, left). These derived metrics are then used for decision support to guide clinicians in diagnosis and treatment.
@@
Conformal prediction (CP) has emerged as a popular, statistically principled uncertainty quantification framework of choice in machine learning, providing guarantees without restrictive distributional assumptions [Vovk et al., 2005, Shafer and Vovk, 2008, Fontana et al., 2023, Papadopoulos et al., 2002, Angelopoulos and Bates, 2021]. While well-studied in the context of typical prediction tasks involving scalar output variables, CP is less explored for tasks such as medical image segmentation, in which the output variables are images. Existing CP methods for segmentation typically focus on deriving bounds for pixel-level errors  [Mossina et al., 2024, Mossina and Friedrich, 2025, Brunekreef et al., 2024, Angelopoulos et al., 2022], which, while useful for understanding variations of local segmentation contours, may yield meaningless or misaligned intervals for downstream derived metrics. On the other hand, a recent study shows that treating the segmentation-to-metric pipeline as a black box and performing CP directly on the output metric space yields intervals that are well-aligned to the metrics (by construction), but are also often inefficient (i.e., large) because the internal biases of the pipeline are not exploited in the vanilla CP formulation [Cheung et al., 2025].
@@
To achieve more efficient intervals, a promising direction is to leverage the powerful inductive biases of neural networks by performing CP on their intermediate representations. Feature Conformal Prediction (FCP) [Teng et al., 2022, Tang et al., 2024, Chen et al., 2024] shows that by working in a semantic feature space, it is possible to generate provably tighter prediction intervals. However, the FCP algorithm requires solving a complex optimization to find the closest adversarial feature vector for each data point, which is computationally prohibitive for high-dimensional feature spaces of typical CNN and transformer architectures which are the workhorses of modern medical image segmentation.
@@
To bridge this gap, we introduce Conformal Metric Perturbation Along Sensitive Subspaces (COMPASS), a framework to perform feature CP in a tractable manner to generate valid and efficient prediction intervals for any (differentiable) metric derived from the output of a neural network. The core concept behind COMPASS is to linearly perturb the features outputted by the network at a particular layer along data-specific directions that are highly sensitive to the metric of interest (Figure 1, center). To make this process tractable for typical neural network layers with a large number of features, we propose computing a low-dimensional manifold of any given layer by applying principal component analysis (PCA) on the gradients of the output metric with respect to each of the layer’s features. This manifold represents the principal directions of sensitivity of the output with respect to that layer. We prove that linear perturbations in the latent space achieve marginal coverage in the output metric space under exchangeability, i.e., for a fresh test point, the ground truth will be contained within the prediction intervals generated by linear feature perturbations. Furthermore, we show that a simple weighted variant of COMPASS may be used to correct for covariate shifts.
@@
We evaluated COMPASS on public datasets for four medical image segmentation tasks: colorectal cancer in histopathology images (EBHI)  [Hu et al., 2023], skin lesion segmentation (HAM10000) [Tschandl et al., 2018], thyroid nodule segmentation in ultrasound images (TN3K) [Gong et al., 2021], and polyp segmentation in endoscopic images (Kvasir) [Jha et al., 2019]. Results show that while standard CP methods achieve valid coverage, they often produce unnecessarily wide prediction intervals. On the other hand, COMPASS finds semantically meaningful directions in the latent space that correspond to monotonic metric changes, resulting in efficient intervals. Furthermore, the weighted extension of COMPASS recovers target coverage under covariate shifts and is the most efficient across all weighted baseline methods. COMPASS paves the way for practical, metric-based uncertainty quantification for medical image segmentation.
Figure 1: Overview of COMPASS. (Left) A medical image segmentation network predicts a segmentation map” from an input image. We conceptually decompose this network into a function $f$ which maps the image to latent features $\hat{z}$, and a function $g$ that maps $\hat{z}$ to the output map. The map may then be used to compute a (differentiable) downstream metric $\hat{y}$ via the function $h$. (Center) We linearly perturb calibration features $\hat{z_{i}}$ in a sample-specific direction $\Delta_{i}$ to find the scores $R_{i}$. The scores are used to find the conformal quantile $\hat{\beta}$. (Right) At test time for subject $n+1$, we perturb the features $\hat{z}_{n+1}$ in the direction $\Delta_{n+1}$ with magnitude $\hat{\beta}$. By Theorem 1, our interval construction is guaranteed to be nested. Therefore, under the assumption of exchangeability, the resulting prediction interval achieves marginal coverage (bottom).
@@ 2.1 Theoretical Coverage under Linear Latent Perturbations
To provide rigorous uncertainty quantification for the metric $Y$, we consider linear perturbations in latent space along a direction $\Delta\in\mathcal{Z}$ and define intervals in the metric space. We first formalize the nestedness condition, which is a fundamental requirement for the validity of any conformal procedure [Vovk et al., 2005, Shafer and Vovk, 2008].
@@
Let $S_{\beta}(x)$ be a family of prediction sets for an input $x\in\mathcal{X}$, parameterized by $\beta\geq 0$. The family $\{S_{\beta}(x)\}_{\beta\geq 0}$ is said to be nested for $x$ if $\beta_{1}\leq\beta_{2}\;\Rightarrow\;S_{\beta_{1}}(x)\subseteq S_{\beta_{2}}(x)$.
This condition guarantees that a larger perturbation magnitude yields a larger (or equal-sized) prediction set, which is necessary for the quantile-based coverage proof. While standard CP methods satisfy this trivially, it becomes a non-trivial condition for deep feature spaces. We now present our main theorem. Crucially, we construct our prediction sets $S_{\beta}(x)$ in a way that guarantees nestedness by definition, thereby ensuring the validity of the conformal procedure.
@@
Let $(X_{i},Y_{i})_{i\geq 1}$ be exchangeable random pairs with $X_{i}\in\mathcal{X}$ and $Y_{i}\in\mathbb{R}$, and split the data into a training set $D_{\mathrm{tr}}$ and a calibration set $D_{\mathrm{cal}}=\{(X_{i},Y_{i})\}_{i=1}^{n}$. Using $D_{\mathrm{tr}}$, fit a segmentation model with decoder $g:\mathcal{Z}\to\mathcal{S}$, and let $\hat{z}(x)\in\mathcal{Z}$ denote the latent vector computed deterministically by the trained encoder for input $x$. Let $h:\mathcal{S}\to\mathbb{R}$ be a measurable metric, and let $\Delta\in\mathcal{Z}$ be any measurable direction that depends only on $D_{\mathrm{tr}}$.
For $x\in\mathcal{X}$ and $\beta\geq 0$, define the metric function along this direction as:
$m_{x}(b):=(h\circ g)(\hat{z}(x)+b\Delta),\quad\text{for }b\in\mathbb{R}.$ (1)
We define the prediction set $S_{\beta}(x)$ as the range of the metric function over the perturbation interval $\beta,+\beta]$:
$S_{\beta}(x):=\left[\min_{b\in[-\beta,+\beta]}m_{x}(b),\max_{b\in[-\beta,+\beta]}m_{x}(b)\right].$ (2)
By construction, this guarantees that the family $\{S_{\beta}(x)\}_{\beta\geq 0}$ is nested (Definition 1). For each calibration pair $(X_{i},Y_{i})$, define the non-conformity score:
$R_{i}:=\inf\{\beta\geq 0:Y_{i}\in S_{\beta}(X_{i})\}\in[0,\infty],$ (3)
@@
$\mathbb{P}\bigl(Y_{n+1}\in S_{\hat{\beta}}(X_{n+1})\;\big|\;D_{\mathrm{tr}}\bigr)\geq 1-\alpha.$ (4)
Proof sketch. Exchangeability of $(X_{i},Y_{i})$ ensures that the rank of the test score $R_{n+1}$ among $\{R_{1},\dots,R_{n},R_{n+1}\}$ is uniformly distributed. By the construction of $S_{\beta}(x)$ as the range $\min(\cdot),\max(\cdot)]$, the nestedness condition is satisfied by definition. Therefore, by construction of $\hat{\beta}$, the standard conformal guarantee holds. See Appendix A for full proof.
@@
Our prediction set $S_{\beta}(x)$ is defined as the metric range over the perturbation interval, which satisfies the nestedness condition required for CP. Our construction forms a conservative envelope that explicitly accounts for any non-monotonic metric behavior within the perturbation interval. Our non-conformity score $R_{i}$ is then the minimal perturbation magnitude $\beta$ required for this envelope to contain the ground truth metric $Y_{i}$.
@@ 2.2 COMPASS: Conformal Metric Perturbation Along Sensitive Subspaces
The proposed framework COMPASS is built on the insight that we can directly calibrate a downstream metric by perturbing a model’s representations along specific, data-dependent directions, $\Delta$. As our theoretical guarantee in Theorem 1 applies to any measurable perturbation direction, the key to an efficient method lies in choosing a direction that is highly sensitive to the metric of interest. We explore two natural choices for this representation space for segmentation: the model’s final output logits and a deeper internal feature layer.
The first and simplest approach (COMPASS-L) operates on the model’s final logits. A uniform, scalar shift applied to this tensor directly modulates the model’s overall confidence before the final activation function. This is equivalent to defining the sensitive direction, $\Delta_{i}$, as a tensor of ones. A richer alternative is to perturb an internal feature representation $\hat{z}$, as done in FCP [Teng et al., 2022]. However, a naive search for an optimal perturbation in high-dimensional spaces is computationally intractable. Furthermore, arbitrary perturbation directions may cause the metric to change in an erratic or non-smooth ways, resulting in unnecessarily wide and inefficient intervals.
To overcome this, we propose a data-driven method (COMPASS-J) to identify a low-dimensional sensitive subspace that is globally effective [^1] Our use of principal directions to restrict the perturbation search space is conceptually related to Belhasin et al. [2023], where authors employed principal directions to construct prediction sets for inverse problems.. Given a metric function in Equation 1 that is differentiable, we first compute Jacobians of the metric $\hat{y}$ with respect to $\hat{z}$ to provide a local, linear map of the metric’s sensitivity with respect to the features for each training sample:
$J_{i}:=\frac{d\,h(g(\hat{z}_{i}))}{d\hat{z}_{i}}\in\mathbb{R}^{C\times D_{1}\times\dots\times D_{3}},$
where $\hat{z}_{i}=f(x_{i})$, $C$ is the number of channels, and $D_{i}$ is the $i$-th spatial dimension. Because the full spatial Jacobian is often too high-dimensional to be practical, we sum the spatial dimensions and apply Principal Component Analysis (PCA) to the set of these vectors from the training set, $\{\mathcal{J}_{i}\}_{i\in D_{\mathrm{tr}}}$, and select the matrix of the top $L$ eigenvectors $V_{L}\in\mathbb{R}^{C\times L}$.
For any given sample, we find its sensitive direction $\mathbf{d}_{i}$ by projecting its sensitivity vector onto this learned subspace, and normalization it to produce direction vector $\Delta_{i}$: